1. Calculating the price elasticity of demand: A step-by-step guide Suppose that during the past year, the price of a laptop computer fell from $2,650 to $2,270. During the same time period, consumer sales increased from 411,000 to 587,000 laptops. Calculate the elasticity of demand between this two price–quantity combinations by using the following steps. After each step, complete the relevant part of the table with the appropriate answers. (Note: For decreases in price or quantity, enter values in the Change column with a minus sign.) Original New Average Change Percentage Change Quantity Price $ $ $ $ Step 1: Fill in the appropriate values for original quantity, new quantity, original price, and new price. Step 2: Calculate the average quantity by adding the original quantity and the new quantity, and then dividing by two. Do the same for the average price. Step 3: Calculate the change in quantity by subtracting the original quantity from the new quantity. Do the same for the change in price. Step 4: Calculate the percentage change in quantity demanded by dividing the change in quantity by the average quantity. Do the same to calculate the percentage change in price. Step 5: Calculate the price elasticity of demand by dividing the percentage change in quantity demanded by the percentage change in price, ignoring the negative sign. Using the midpoint method, the elasticity of demand for laptops is about.
Answer #1
Original, O New, N Average, A =(O+N)/2 Change, D=N-O Percentage Change= D/A
Quantity 411000 587000 499000 176000 35.2705%
Price $ 2650 2270 2460 -380 -15.4472%

Step 1

We can directly plug in Original and New values of price and quantity (as these are given) as shown in above table.

Step 2

Average Quantity=(New+Original)/2=(411000+587000)/2=499000

Average Quantity=(New+Original)/2=(2270+2650)/2=$2460

Step 3

Change in quantity=New-Original =587000-411000=176000

Change in price=New-Original =2270-2650=-$380

Step 4

Percentage Change in quantity=Change in quantity/Average quantity=176000/499000=35.2705%

Percentage Change in price=Change in price/Average price=-380/2460=-15.4472%

Step 5

Ignoring the negative sign in change in price,

Price elasticity of demand=Percentage change in quantity/Percentage change in price =35.2705%/15.4472%=2.28

Using the midpoint method, the elasticity of demand for laptops is about 2.28